# Nonlinear Bayesian Filtering with Natural Gradient Gaussian Approximation

### v1

@@ Abstract
Practical Bayes filters often assume the state distribution of each time step to be Gaussian for computational tractability, resulting in the so-called Gaussian filters. When facing nonlinear systems, Gaussian filters such as extended Kalman filter (EKF) or unscented Kalman filter (UKF) typically rely on certain linearization techniques, which can introduce large estimation errors. To address this issue, this paper reconstructs the prediction and update steps of Gaussian filtering as solutions to two distinct optimization problems, whose optimal conditions are found to have analytical forms from Stein’s lemma. It is observed that the stationary point for the prediction step requires calculating the first two moments of the prior distribution, which is equivalent to that step in existing moment-matching filters. In the update step, instead of linearizing the model to approximate the stationary points, we propose an iterative approach to directly minimize the update step’s objective to avoid linearization errors. For the purpose of performing the steepest descent on the Gaussian manifold, we derive its natural gradient that leverages Fisher information matrix to adjust the gradient direction, accounting for the curvature of the parameter space. Combining this update step with moment matching in the prediction step, we introduce a new iterative filter for nonlinear systems called Natural Gradient Gaussian Approximation filter, or NANO filter for short. We prove that NANO filter locally converges to the optimal Gaussian approximation at each time step. The estimation error is proven exponentially bounded for nearly linear measurement equation and low noise levels through contructing a supermartingale-like inequality across consecutive time steps. Real-world experiments demonstrate that, compared to popular Gaussian filters such as EKF, UKF, iterated EKF, and posterior linearization filter, NANO filter reduces the average root mean square error by approximately 45% while maintaining a comparable computational burden.
@@ I Introduction
It turns out that the design philosophy of Gaussian filters generally consists of two steps: (i) approximate the nonlinear models to linear forms with additive Gaussian noise, and (ii) perform KF based on this linear Gaussian model. Within this framework, the differences between Gaussian filters are primarily attributed to different linearization techniques. The earliest technique involves directly using the Taylor series approximation to linearize the nonlinear function. Typically, the first-order expansion is employed to avoid the occurrence of tensors in high-order expansions, exemplified by the well-known extended Kalman filter (EKF) [6, 7]. Originally developed by NASA for navigation tasks, EKF linearizes the nonlinear function around the state estimate and performs KF using the resulting affine-form system equation. Building on this foundation, iterated extended Kalman filter (IEKF) [8] repeatedly performs linearization at each updated approximation of the posterior mean instead of the prior mean as in EKF. In fact, the iterative process of this method is a Gauss-Newton iteration that essentially linearizes the model around the maximum a posteriori (MAP) estimate of the state [9].
@@
  • We prove that the NANO filter locally converges to the optimal Gaussian approximation at each time step, with accuracy up to a second-order remainder in the Taylor expansion. Additionally, the estimation error is proven exponentially bounded for nearly linear measurement equation and low noise levels through contructing a supermartingale-like inequality across consecutive time steps. We further show that NANO filter can naturally extend the Bayesian posterior to Gibbs posterior, enabling the use of more flexible loss functions to enhance robustness against outliers in measurement data. On its basis, three robust variants of NANO filter are introduced, each employing different robust loss function designs. Simulations and real-world experiments demonstrate that NANO filter and its robust variants significantly outperforms popular Gaussian filters, such as EKF, UKF, IEKF, and PLF for nonlinear systems.
The remainder of this paper is structured as follows: Section II formulates the problem, and Section III explores the optimal Gaussian approximation. Section IV introduces the NANO filter algorithm, followed by theoretical analysis in Section V. Section VI proposes the robust variants of NANO filter. Section VII provides a discussion on the proposed algorithm. Simulations and real-world experiments are presented in Section VIII.
Notation: All vectors are considered as column vectors. The symbol $D_{\mathrm{KL}}(p\|q)=\mathbb{E}_{p}\left\{\log{\frac{p}{q}}\right\}$ denotes the KL divergence between two probability distributions $p$ and $q$. Unless otherwise specified, $\|x\|$ refers to the $\ell_{2}$-norm of the vector $x$, defined as $\|x\|=\sqrt{x^{\top}x}$. The notation $\mathcal{N}(x;\mu,\Sigma)$ represents the Gaussian probability density function for the variable $x$ with mean $\mu$ and covariance matrix $\Sigma$. For simplicity, this may be abbreviated as $\mathcal{N}(\mu,\Sigma)$. Furthermore, $\mathbb{E}\{x\}$, $\mathbb{D}\{x\}$, and $\mathrm{Cov}(x,y)$ denote the expectation, variance, and covariance, respectively. The notation $A\leq B$ ($A<B$) for symmetric matrices $A$ and $B$ indicates that $B-A$ is positive (semi-)definite. The symbol $\otimes$ denotes the Kronecker product.
@@ II Problem Statement
where $x_{t}\in\mathbb{R}^{n}$ is the system state, $y_{t}\in\mathbb{R}^{m}$ is the noisy measurement. The function ${f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is referred to as the transition function, while ${g}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is called measurement function; ${\xi}_{t}$ denotes process noise, and ${\zeta}_{t}$ represents measurement noise. Typically, the initial state $x_{0}$, the process noise $\{\xi_{t}\}$, and the measurement noise $\{\zeta_{t}\}$ are all mutually independent, with the process and measurement noises being independent and identically distributed across time. This state-space model description in (1) can be transformed into a hidden Markov model:
@@
Here, $p(x_{t}|x_{t-1})$ and $p(y_{t}|x_{t})$ represent the transition and output probabilities respectively while $p(x_{0})$ denotes the initial state distribution. In essence, (2) and (1) are different representations of the same system model. For example, consider the transition model $x_{t}=Ax_{t-1}+\xi_{t-1}$, where $\xi_{t-1}\sim\mathcal{N}(\xi_{t-1};0,Q)$, with $Q$ denoting the covariance matrix of the process noise. This can be equivalently described by the transition probability $p(x_{t}|x_{t-1})=\mathcal{N}(x_{t};Ax_{t-1},Q)$.
@@
Linearization Technique Basic Principle $\mathcal{N}(y;g(x),\Sigma)\approx\mathcal{N}(y;Ax+b,\Lambda)$ Representative Algorithms
@@
Statistical Linear Regression $\operatorname*{arg\,min}_{A,b}\mathbb{E}_{x}\left\{(y-Ax-b)(y-Ax-b)^{\top}\right\}$Downarrow $A=\bm{\mathrm{Cov}}(x,y)^{\top}\mathbb{D}(x),b=\mathbb{E}\left\{y\right\}-A\mathbb{E}\left\{x\right\},\Lambda=\mathbb{D}\left\{y\right\}-A\mathbb{D}\left\{x\right\}A^{\top}$ UKF [11], Gauss–Hermite KF [12], cubature KF [13], and PLF[14]
@@ III-A Optimization Viewpoint of Bayesian Filtering
Note that in both (5) and (6), $q:\mathbb{R}^{n}\to\mathbb{R}$ represents the candidate density function. Besides, we use the notation $\mathbb{E}_{\begin{subarray}{c}p(x)\\ p(y)\end{subarray}}\left\{f(x,y)\right\}$ to represent the expectation of $f(x,y)$ with respect to both distributions $p(x)$ and $p(y)$, i.e., $\mathbb{E}_{\begin{subarray}{c}p(x)\\ p(y)\end{subarray}}\left\{f(x,y)\right\}:=\mathbb{E}_{p(x)}\mathbb{E}_{p(y)}\left\{f(x,y)\right\}$.
@@
If we choose $Z=\int p(x_{t}|x_{t-1})p(x_{t-1}|y_{1:t-1})\,\mathrm{d}x_{t-1}$ and recall that the Kullback-Leibler (KL) divergence reaches its minimum value uniquely when its arguments are identical, we find that the solution to (7) is indeed
@@ III-B Structure of Stationary Points of $L(\hat{x}_{t},P_{t})$
$\Sigma^{*}$ $=\mathbb{E}_{p(x)}\left\{(x-\mu^{*})(x-\mu^{*})^{\top}\right\}$ (11b)
$=\mathbb{E}_{p(x)}\left\{xx^{\top}\right\}-\mu^{*}{\mu^{*}}^{\top}.$ (11c)
@@ III-C Structure of Stationary Points of $J(\hat{x}_{t},P_{t})$
$\frac{\partial{J}(\hat{x}_{t|t},P_{t|t})}{\partial{\hat{x}_{t|t}}}=\frac{\partial{J}(\hat{x}_{t|t},P_{t|t})}{\partial{P_{t|t}}}=0,$
@@ IV Natural Gradient Gasussian Approximation
In the previous section, we observed that existing moment-matching KF methods already provide approximate solutions for the optimal Gaussian approximation in the prediction step, but the update step remains inadequately addressed in existing Gaussian filters. In this section, we aim to tackle Q2. Specifically, we will design an algorithm to solve (9b).
By examining the structure of the extreme conditons defined by (18), we find that it is challenging to directly obtain an analytical form of the stationary point because it is typically impossible to isolate the updates for the mean and covariance in (18). Therefore, a more practical approach to finding the optimal solution is to directly minimize the update cost $J(\hat{x}_{t},P_{t})$. To find the steepest descent in optimizing the parameters of Gaussian distributions [21, 22], we derive an natural gradient iteration for finding the optimal Gaussian approximation.
@@
$\left(P_{t}^{-1}\right)^{(i+1)}=$ $\left(P_{t}^{-1}\right)^{(i)}$
@@
To address the first issue, we could use well-established numerical integration methods, such as the unscented transform [11], Gauss–Hermite integration [26], or spherical cubature integration [13] to approximate the expectation calculations. For the second issue, one possible and efficient solution is to provide a sufficiently good initialization. For example, we could solve the MAP estimation problem:
$\hat{x}_{t|t}^{\mathrm{MAP}}$ $=\operatorname*{arg\,max}_{x_{t}}\left\{\mathcal{N}(x_{t};\hat{x}_{t|t-1},P_{t|t-1})\cdot\exp\{\ell(x_{t},y_{t})\}\right\}$ (29)
$=\operatorname*{arg\,max}_{x_{t}}\left\{\log{\mathcal{N}(x_{t};\hat{x}_{t|t-1},P_{t|t-1})}+\ell(x_{t},y_{t})\right\},$
and use Laplace’s approximation [27] to construct the initial mean and covariance in (28), as shown in the subsequent equation:
@@
$\left(P_{t}^{-1}\right)^{(0)}=$ $\left.\frac{\partial^{2}\left\{-\log{\mathcal{N}(x_{t};\hat{x}_{t|t-1},P_{t|t-1})}-\ell(x_{t},y_{t})\right\}}{\partial x_{t}^{2}}\right|_{x_{t}=\hat{x}_{t|t}^{\mathrm{MAP}}}.$
@@
An important consideration for the iterative scheme is the stopping criterion. We propose using the KL divergence between two consecutive Gaussian distributions, specifically $\mathcal{N}^{(i)}=\mathcal{N}(x_{t};\hat{x}_{t|t}^{(i)},P_{t|t}^{(i)})$ and $\mathcal{N}^{(i+1)}=\mathcal{N}(x_{t};\hat{x}_{t|t}^{(i+1)},P_{t|t}^{(i+1)})$ to determine when to stop the iteration. As suggested by [14], the stopping condition can be defined as
@@
where $\gamma$ is a predefined threshold. This approach is more effective compared to using $D_{\mathrm{KL}}(\mathcal{N}^{(i+1)}\|\mathcal{N}^{(i)})<\gamma$ because $\mathcal{N}^{(i+1)}$ is generally more concentrated. By using $D_{\mathrm{KL}}(\mathcal{N}^{(i)}\|\mathcal{N}^{(i+1)})$, we better capture meaningful changes in the iteration and ensure that the criterion is sensitive enough to detect convergence, while avoiding premature termination of the algorithm.
@@ V Theoretical Analysis
In this section, we will answer Q3, i.e., we will provide the stability and convergence analysis for NANO filter.
@@ V-A Convergence Analysis
This theorem indicates that the natural gradient descent iteration in the update step of the NANO filter provides a guarantee of local convergence. This guarantee is achieved by approximating the objective function with second-order accuracy around the mean and first-order accuracy around the inverse of the covariance matrix.
The proof of Theorem 1 is provided in Appendix -B. The key idea is to show that the difference in update costs between consecutive iterations can be expressed as a semi-negative definite quadratic form. Technically, achieving this semi-negative definiteness relies on the use of the Fisher information matrix, which corrects the gradient direction to provide the steepest descent on the Gaussian manifold. This adjustment makes the gradient ”natural” because it aligns with the geometry of the Riemannian space of Gaussian parameters.
For linear Gaussian systems, the NANO filter achieves the optimal solution of equation (9b) in a single iteration, as stated in the following corollary:

### v2

@@ Abstract
Practical Bayes filters often assume the state distribution of each time step to be Gaussian for computational tractability, resulting in the so-called Gaussian filters. When facing nonlinear systems, Gaussian filters such as extended Kalman filter (EKF) or unscented Kalman filter (UKF) typically rely on certain linearization techniques, which can introduce large estimation errors. To address this issue, this paper reconstructs the prediction and update steps of Gaussian filtering as solutions to two distinct optimization problems, whose optimal conditions are found to have analytical forms from Stein’s lemma. It is observed that the stationary point for the prediction step requires calculating the first two moments of the prior distribution, which is equivalent to that step in existing moment-matching filters. In the update step, instead of linearizing the model to approximate the stationary points, we propose an iterative approach to directly minimize the update step’s objective to avoid linearization errors. For the purpose of performing the steepest descent on the Gaussian manifold, we derive its natural gradient that leverages Fisher information matrix to adjust the gradient direction, accounting for the curvature of the parameter space. Combining this update step with moment matching in the prediction step, we introduce a new iterative filter for nonlinear systems called Natural Gradient Gaussian Approximation filter, or NANO filter for short. We prove that NANO filter locally converges to the optimal Gaussian approximation at each time step. Furthermore, the estimation error is proven exponentially bounded for nearly linear measurement equation and low noise levels through constructing a supermartingale-like property across consecutive time steps. Real-world experiments demonstrate that, compared to popular Gaussian filters such as EKF, UKF, iterated EKF, and posterior linearization filter, NANO filter reduces the average root mean square error by approximately 45% while maintaining a comparable computational burden.
@@ I Introduction
It turns out that the design philosophy of Gaussian filters generally consists of two steps: (i) approximate the nonlinear models to linear forms with additive Gaussian noise, and (ii) perform KF based on this linear Gaussian model. Within this framework, the differences between Gaussian filters are primarily attributed to different linearization techniques. The earliest technique involves directly using the Taylor series approximation to linearize the nonlinear function. Typically, the first-order expansion is employed to avoid the occurrence of tensors in high-order expansions, exemplified by the well-known extended Kalman filter (EKF) [6, 7]. Originally developed by NASA for navigation tasks, EKF linearizes the nonlinear function around the state estimate and performs KF using the resulting affine-form system equation. Building on this foundation, iterated extended Kalman filter (IEKF) [8] repeatedly performs linearization at each updated approximation of the posterior mean instead of the prior mean as in EKF. In fact, the iterative process of this method is a Gauss-Newton iteration that essentially linearizes the model around the maximum a posteriori estimate of the state [9].
@@
  • We prove that the NANO filter locally converges to the optimal Gaussian approximation at each time step, with accuracy up to a second-order remainder in the Taylor expansion. Additionally, the estimation error is proven exponentially bounded for nearly linear measurement equation and low noise levels through contructing a supermartingale-like property across consecutive time steps. We further show that NANO filter can naturally extend the Bayesian posterior to Gibbs posterior, enabling the use of more flexible loss functions to enhance robustness against outliers in measurement data. On its basis, three robust variants of NANO filter are introduced, each employing different robust loss function designs. Simulations and real-world experiments demonstrate that NANO filter and its robust variants significantly outperforms popular Gaussian filters, such as EKF, UKF, IEKF, and PLF for nonlinear systems.
The remainder of this paper is structured as follows: Section II formulates the problem, and Section III explores the optimal Gaussian approximation. Section IV introduces the NANO filter algorithm, followed by theoretical analysis in Section V. Section VI proposes the robust variants of NANO filter. Section VII provides a discussion on the proposed algorithm. Simulations and experiments are presented in Section VIII.
Notation: All vectors are considered as column vectors. The symbol $D_{\mathrm{KL}}(p\|q)=\mathbb{E}_{p}\left\{\log{\frac{p}{q}}\right\}$ denotes the Kullback-Leibler (KL) divergence between two probability distributions $p$ and $q$. Unless otherwise specified, $\|x\|$ refers to the $\ell_{2}$-norm of the vector $x$, defined as $\|x\|=\sqrt{x^{\top}x}$. The notation $\mathcal{N}(x;\mu,\Sigma)$ represents the Gaussian probability density function for the variable $x$ with mean $\mu$ and covariance matrix $\Sigma$. For simplicity, this may be abbreviated as $\mathcal{N}(\mu,\Sigma)$. Furthermore, $\mathbb{E}\{x\}$, $\mathbb{D}\{x\}$, and $\mathrm{Cov}(x,y)$ denote the expectation, variance, and covariance, respectively. The notation $A\leq B$ ($A<B$) for symmetric matrices $A$ and $B$ indicates that $B-A$ is positive (semi-)definite. The symbol $\otimes$ denotes the Kronecker product.
@@ II Problem Statement
where $x_{t}\in\mathbb{R}^{n}$ is the system state, $y_{t}\in\mathbb{R}^{m}$ is the noisy measurement. The function ${f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is referred to as the transition function, while ${g}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is called measurement function; ${\xi}_{t}$ denotes process noise, and ${\zeta}_{t}$ represents measurement noise. Typically, the initial state $x_{0}$, the process noise $\{\xi_{t}\}$, and the measurement noise $\{\zeta_{t}\}$ are all mutually independent, with the process and measurement noises being independent and identically distributed across time. This state-space model description in (1) can be represented as a hidden Markov model:
@@
Here, $p(x_{t}|x_{t-1})$ and $p(y_{t}|x_{t})$ are the transition and output probabilities respectively while $p(x_{0})$ denotes the initial state distribution. In essence, (2) and (1) are different representations of the same system model. For example, consider the transition model $x_{t}=Ax_{t-1}+\xi_{t-1}$, where $\xi_{t-1}\sim\mathcal{N}(\xi_{t-1};0,Q)$, with $Q$ denoting the covariance matrix of the process noise. This can be equivalently described by the transition probability $p(x_{t}|x_{t-1})=\mathcal{N}(x_{t};Ax_{t-1},Q)$.
@@
Linearization Technique Basic Principle: $\mathcal{N}(y;g(x),\Sigma)\approx\mathcal{N}(y;Ax+b,\Lambda)$ Representative Algorithms
@@
Statistical Linear Regression $\operatorname*{arg\,min}_{A,b}\mathbb{E}_{x}\left\{(y-Ax-b)^{\top}(y-Ax-b)\right\}$Downarrow $A=\bm{\mathrm{Cov}}(x,y)^{\top}\mathbb{D}(x),b=\mathbb{E}\left\{y\right\}-A\mathbb{E}\left\{x\right\},\Lambda=\mathbb{D}\left\{y\right\}-A\mathbb{D}\left\{x\right\}A^{\top}$ UKF [11], Gauss–Hermite KF [12], cubature KF [13], and PLF[14]
@@ III-A Optimization Viewpoint of Bayesian Filtering
Note that in both (5) and (6), $q:\mathbb{R}^{n}\to\mathbb{R}$ represents the candidate density function. Besides, we use the notation $\mathbb{E}_{\begin{subarray}{c}p(x)\\ p(y)\end{subarray}}\left\{f(x,y)\right\}$ to represent the expectation of $f(x,y)$ with respect to both distributions $p(x)$ and $p(y)$, i.e., $\mathbb{E}_{\begin{subarray}{c}p(x)\\ p(y)\end{subarray}}\left\{f(x,y)\right\}\triangleq\mathbb{E}_{p(x)}\mathbb{E}_{p(y)}\left\{f(x,y)\right\}$.
@@
If we choose $Z=\int p(x_{t}|x_{t-1})p(x_{t-1}|y_{1:t-1})\,\mathrm{d}x_{t-1}$ and recall that the KL divergence reaches its minimum value uniquely when its arguments are identical, we find that the solution to (7) is indeed
@@ III-B Structure of Stationary Points of $L(\hat{x}_{t},P_{t})$
$\Sigma^{*}$ $=\mathbb{E}_{p(x)}\left\{(x-\mu^{*})(x-\mu^{*})^{\top}\right\}$
$=\mathbb{E}_{p(x)}\left\{xx^{\top}\right\}-\mu^{*}{\mu^{*}}^{\top}.$ (11b)
@@ III-C Structure of Stationary Points of $J(\hat{x}_{t},P_{t})$
$\frac{\partial{J}(\hat{x}_{t|t},P_{t|t})}{\partial{\hat{x}_{t|t}}}=\frac{\partial{J}(\hat{x}_{t|t},P_{t|t})}{\partial{P_{t|t}^{-1}}}=0,$
@@ IV Natural Gradient Gasussian Approximation
In the previous section, we observed that existing moment-matching KF methods already provide exact numerical solutions for the optimal Gaussian approximation in the prediction step. However, the update step is not sufficiently resolved in current Gaussian filters, as these methods rely on linearization to approximate the stationary point, which introduce linearization errors. In this section, we aim to tackle Q2. Specifically, we will design an algorithm to solve (9b).
By examining the structure of the extreme conditons defined by (18), we find that it is challenging to directly obtain an analytical form of the stationary point because it is typically impossible to isolate the updates for the mean and covariance in (18). Therefore, a more practical approach to finding the optimal solution is to directly minimize the update cost $J(\hat{x}_{t},P_{t})$. To find the steepest descent in optimizing the parameters of Gaussian distributions [21, 22], we derive a natural gradient iteration for finding the optimal Gaussian approximation.
@@
$\left(P_{t}^{-1}\right)^{(i+1)}=$ $\left(P_{t}^{-1}\right)^{(i)}$ (25)
@@
To address the first issue, we could use well-established numerical integration methods, such as the unscented transform [11], Gauss–Hermite integration [26], or spherical cubature integration [13] to approximate the expectation calculations. For the second issue, one possible and efficient solution is to provide a sufficiently good initialization. For example, we could solve the maximum a posterior estimation problem:
$\hat{x}_{t|t}^{\mathrm{MAP}}$ $=\operatorname*{arg\,max}_{x_{t}}\left\{\mathcal{N}(x_{t};\hat{x}_{t|t-1},P_{t|t-1})\cdot\exp\{-\ell(x_{t},y_{t})\}\right\}$ (29)
$=\operatorname*{arg\,max}_{x_{t}}\left\{\log{\mathcal{N}(x_{t};\hat{x}_{t|t-1},P_{t|t-1})}-\ell(x_{t},y_{t})\right\},$
and use Laplace’s approximation [27] to construct the initial mean and covariance for the iteration in (28), as shown in the subsequent equation:
@@
$\left(P_{t}^{-1}\right)^{(0)}=$ $\left.\frac{\partial^{2}\left\{-\log{\mathcal{N}(x_{t};\hat{x}_{t|t-1},P_{t|t-1})}+\ell(x_{t},y_{t})\right\}}{\partial x_{t}^{2}}\right|_{x_{t}=\hat{x}_{t|t}^{\mathrm{MAP}}}.$
@@
Another important consideration for the iterative scheme is the stopping criterion. As suggested by [14], we use the KL divergence between two consecutive Gaussian distributions, specifically $\mathcal{N}^{(i)}=\mathcal{N}(x_{t};\hat{x}_{t|t}^{(i)},P_{t|t}^{(i)})$ and $\mathcal{N}^{(i+1)}=\mathcal{N}(x_{t};\hat{x}_{t|t}^{(i+1)},P_{t|t}^{(i+1)})$ to determine when to stop the iteration:
@@
where $\gamma$ is a predefined threshold. This approach is more effective compared to using $D_{\mathrm{KL}}(\mathcal{N}^{(i+1)}\|\mathcal{N}^{(i)})<\gamma$ because $\mathcal{N}^{(i+1)}$ is generally more concentrated. By using $D_{\mathrm{KL}}(\mathcal{N}^{(i)}\|\mathcal{N}^{(i+1)})$, we ensure that the criterion remains sensitive to convergence while avoiding premature termination of the algorithm.
@@ V Theoretical Analysis
In this section, we will answer Q3, i.e., we will provide the convergence and stability analysis for NANO filter.
@@ V-A Convergence Analysis
This theorem indicates that natural gradient descent iteration in the update step of the NANO filter provides a guarantee of local convergence. This guarantee is achieved by approximating the objective function with second-order accuracy around the mean and first-order accuracy around the inverse of the covariance matrix.
The proof of Theorem 1 is provided in Appendix -B. The key idea is to show that the difference in update costs between consecutive iterations can be expressed as a semi-negative definite quadratic form. Technically, achieving this semi-negative definiteness relies on the use of the Fisher information matrix, which corrects the gradient direction to provide the steepest descent on the Gaussian manifold. This adjustment makes the gradient “natural” because it aligns with the geometry of the Riemannian space of Gaussian parameters.
For linear Gaussian systems, the NANO filter achieves the optimal solution of (9b) in a single iteration, as stated in the following corollary: